OpenStudy (anonymous):

27^-4/3 equals what?

OpenStudy (mos1635):

1/81

OpenStudy (anonymous):

thanks so much :)

OpenStudy (callisto):

(27)^(-4/3) =(3^3)^(-4/3) =3^[3(-4/3)] = 3^-4 = 1/ 3^4 = 1/81

OpenStudy (anonymous):

\[27^{-\frac{4}{3}} = (27^{\frac{1}{3}})^{-4} = 3^{-4} = \frac{1}{3^4} = \frac{1}{81}\]

OpenStudy (mos1635):

(27^1/3)^(-4) 3^(-4) 1/81

OpenStudy (cwrw238):

= 1 divided by ( cube root of 27 ) to the power 4

OpenStudy (anonymous):

medals for everyone :D

OpenStudy (anonymous):

OMG I got into Harvard University. Just checked my mail!!!!!!!!!!!!!

OpenStudy (anonymous):

awesome! well done! when do you go?

OpenStudy (callisto):

Congrats!

OpenStudy (anonymous):

i have to start in summer since am taking extra classes

OpenStudy (anonymous):

and what are you going to do there?

OpenStudy (anonymous):

i will be top in the school and then go to harvard medical school

OpenStudy (anonymous):

since i'm val dictorian in my school right now, i shouldn't have a problem. even cosidering i go to Pingry.

OpenStudy (anonymous):

im going to do maths at uni next year

OpenStudy (anonymous):

at what??

OpenStudy (anonymous):

university? im from the uk, is it called college over there?

OpenStudy (anonymous):

oh yeah it is

OpenStudy (anonymous):

you must be uber smart to get harvard , congrats

OpenStudy (anonymous):

thanks, but it is expected of me so i don't understand the fuss. i'm indian. i even have a job all set. One of my dad's clients so my potential so in Harvard, i'll have some extra money

OpenStudy (anonymous):

you certainly have the confidence :P

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

once you're president can you make it so everyone get free ice cream?